OpenStudy (anonymous):

what is an equation for a line that passes through (-9,6) and (9,-6)

OpenStudy (anonymous):

No it is only gives me (-9,6) and (9,-6)

OpenStudy (anonymous):

so its 6-(-6) over (-9)-9 so 12 over 18 or 0 over 18

OpenStudy (anonymous):

no response...

OpenStudy (anonymous):

okay so 2 over 3

OpenStudy (anonymous):

what now

OpenStudy (anonymous):

can't you do it for me <3 XD

OpenStudy (anonymous):

what do they stalk people and ban them immediately for giving an answer

OpenStudy (anonymous):

the day i get suspended from a website that is used for school help is the day obama gets called a good president

OpenStudy (anonymous):

lmao okay

OpenStudy (anonymous):

i just googled it and it said y=-0.67x

OpenStudy (anonymous):

lmao i don't want to sit here and try to solve these

OpenStudy (anonymous):

but was it right

OpenStudy (anonymous):

are you writing a novel like what are you doing

OpenStudy (anonymous):

texaschic101 has been writing for atleast 2 minutes

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (-9,6)...x1 = -9 and y1 = 6 (9,-6)...x2 = 9 and y2 = -6 now we sub slope(m) = (-6 - 6) / (9 - (-9) slope(m) = -12 / (9 + 9) slope(m) = -12/18 which reduces to - 2/3 now use y = mx + b slope(m) = -2/3 use either of your points...(-9,6)...x = -9 and y = 6 now sub y = mx + b 6 = -2/3(-9) + b 6 = 18/3 + b 6 = 6 + b 6 - 6 = b 0 = b so your equation is : y = -2/3x + 0....or just y = -2/3x